OpenStudy (anonymous):

Please help!!! A scale drawing of a storage box on a coordinate grid shows a square with the points (-3, 5) and (1, 7) as the locations of two adjacent vertices. Which pair of coordinate points could represent the other two vertices? (0, 1) and (3, 3) (-5, 9) and (1, 11) (-5, 9) and (-1, 11) (0, 1) and (4, 3)

OpenStudy (anonymous):

@warningiflirt can you help

OpenStudy (anonymous):

@love_jessika15 can you help please

OpenStudy (anonymous):

you need to Find how far you move in x, and in y to get from the starting point to the ending point (the 2 they gave us) start at (-3,5) add 4 to the x, add 2 to the y -3+4, 5+2 => 1,7 which is the 2nd point Now do it with each of the possible answers: (-5, 9) and (-1, 11) -5+4, 9+2= -1,11 (it matches!) (0, 1) and (4, 3) also matches, so instead we use the distance formula, for the original pairs d^2 = (1 - -3)^2 + (7-5)^2 = 4^2 + 2^2 = 16+4= 20 (-3-(-5))^2+(9-5)^2=4+16=20 for pair -5,9 answer (-5, 9) and (-1, 11)

OpenStudy (anonymous):

alright one sec

OpenStudy (anonymous):

if that made any sense to you lol

OpenStudy (anonymous):

@warningiflirt ok I understand, wow thanks alot

OpenStudy (anonymous):

yw